How many ounces of a 50 % acetic acid solution should a photographer add to 32 ounces of a 5 % solution to obtain 30 % acetic acid
Answer by mananth(16949)

-------------- percent ---------------- Amount
Acetic acid I 5 ---------------- 32 Ounces
Acetic acid II 50 ---------------- x Ounces
Mixture 30 ---------------- 32 + x Ounces

5*32+50 x =30(32+x)
160+50x =960+30x
50x-30x =-160+ 960
20x=800
/ 20
x = 40 Ounces Acetic acid II

m.ananth@hotmail.ca